OpenStudy (anonymous):

use this information to solve each problem answers have to be in simplest form. alexandra doubles a recipe for punch that needs 3 ingredients 6 and 2 thirds cups of o.j. 3 and 1 half cups of cranberry juice , and 4 and 3 fourths cups of ginger ale. how many cups of punch is ally making show all of your work and write your answer as a complete sentence

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

who are you?

OpenStudy (anonymous):

wow ur mom must b so proud

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

hey michele

OpenStudy (michele_laino):

Hi! @leovaldez

OpenStudy (anonymous):

ur awesome 4 helping me

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

sooo

OpenStudy (anonymous):

help?

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (anonymous):

kk take ur time

OpenStudy (anonymous):

so done pondering?

OpenStudy (michele_laino):

we have these quantities: 6 2/3 = 20/3 cups; 3 1/2= 7/2 cups; 4 3/4 = 19/4 cups, so in total we have: \[\frac{{20}}{3} + \frac{7}{2} + \frac{{19}}{4} = ...?\] plese complete that computation

OpenStudy (michele_laino):

@leovaldez

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

and how do u simplify it?

OpenStudy (michele_laino):

what is the least common multiple of 3, 2 and 4?

OpenStudy (anonymous):

? idk

OpenStudy (michele_laino):

hint: \[\begin{gathered} 4 = {2^2} \times 1 \hfill \\ 3 = 3 \times 1 \hfill \\ 2 = 2 \times 1 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

you have to pick all common and uncommon factors with the highest exponent

OpenStudy (anonymous):

what is the ans im tired i can't think

OpenStudy (anonymous):

i need 2 sleep but not b 4 this work is done i have 2 more

OpenStudy (michele_laino):

I'm sorry, I can not give the direct answer, since I have to respect the Code Of Conduct

OpenStudy (anonymous):

oh im sorry i wasn't thinking see im so tired

OpenStudy (usukidoll):

\[\frac{{20}}{3} + \frac{7}{2} + \frac{{19}}{4} = ...? \] we can't add these fractions until they have a common denominator.

OpenStudy (anonymous):

okkkkk

OpenStudy (anonymous):

i did it

OpenStudy (anonymous):

so i don't get it

OpenStudy (usukidoll):

huh?

OpenStudy (anonymous):

i added them but it needs 2 b simplifyed

OpenStudy (michele_laino):

what is your result? @leovaldez

OpenStudy (anonymous):

46 over 9

OpenStudy (usukidoll):

if you're just adding numerators without finding the least common denominator....I don't believe this.

OpenStudy (anonymous):

wat?

OpenStudy (usukidoll):

I know where you got that 9 from 9 = 3+2+4 I said you can't add these fractions because they have different denominators. All of the denominators have to be the same

OpenStudy (anonymous):

so how do i do that?

OpenStudy (usukidoll):

find the least common denominator.. or you can just grab all the denominators and multiply them out.

OpenStudy (usukidoll):

and then figure out what two numbers are missing for each fraction .

OpenStudy (anonymous):

i know i sound dumb but hey im super tired it's 11:20 where i amggggggggggggggggg lyufhhhhhhhhhhhhkefhyxfyhfgeyx.fylefyglf

OpenStudy (usukidoll):

and I don't want to get in trouble in front of a qualified helper.

OpenStudy (anonymous):

sorry im awake im awake!

OpenStudy (michele_laino):

@leovaldez please complete these computations: \[\begin{gathered} \frac{{20}}{3} = \frac{{20 \times 4}}{{3 \times 4}} = ... \hfill \\ \hfill \\ \frac{7}{2} = \frac{{7 \times 6}}{{2 \times 6}} = ... \hfill \\ \hfill \\ \frac{{19}}{4} = \frac{{19 \times 3}}{{4 \times 3}} = ... \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there

OpenStudy (anonymous):

did it

OpenStudy (michele_laino):

what did you get?

OpenStudy (michele_laino):

@leovaldez